# Towards Explaining Deep Neural Network Compression Through a Probabilistic Latent Space

### v1

@@ Abstract
Despite the impressive performance of deep neural networks (DNNs), their computational complexity and storage space consumption have led to the concept of network compression. While DNN compression techniques such as pruning and low-rank decomposition have been extensively studied, there has been insufficient attention paid to their theoretical explanation. In this paper, we propose a novel theoretical framework that leverages a probabilistic latent space of DNN weights and explains the optimal network sparsity by using the information-theoretic divergence measures. We introduce new analogous projected patterns (AP2) and analogous-in-probability projected patterns (AP3) notions for DNNs and prove that there exists a relationship between AP3/AP2 property of layers in the network and its performance. Further, we provide a theoretical analysis that explains the training process of the compressed network. The theoretical results are empirically validated through experiments conducted on standard pre-trained benchmarks, including AlexNet, ResNet50, and VGG16, using CIFAR10 and CIFAR100 datasets. Through our experiments, we highlight the relationship of AP3 and AP2 properties with fine-tuning pruned DNNs and sparsity levels.
1 Introduction
@@
A wide variety of real-world problems are investigated by deep neural networks (DNNs) which enable powerful learning across a wide range of task learning. DNNs have provided remarkable performance in solving complex problems and this has led to significant growth of interest in the area of neural networks. However; the higher the performance, the more the number of parameters, resulting in complexity and space consumption. While DNNs pruning methods [1, 2, 3, 4, 5, 6, 7, 8] compress networks and tackle computational complexity challenges, due to the lack of sufficient theoretical explanation of such methods, it’s difficult to understand their behaviors. In this paper, we propose a novel framework that explains DNN pruning using probabilistic latent spaces of DNN weights and divergence measures. Our main idea is built upon mapping the weight matrix of the layers in both the original network and its corresponding sparse/pruned network to probability spaces from a distribution where the weight matrices are the parameter sets. To clarify our approach: consider a network $F^{(L)}$ with $L$-layers and its sparse version, denoted by $\widetilde{F}^{(L)}$ with parameters $\omega$ and $\widetilde{\omega}=m\odot\omega$, respectively, in which $m$ is a binary mask. Suppose that $f^{(l)}$, filters of the $l$-th layer of $F^{(L)}$ with $M_{l}$ neurons has weight matrix of $\omega^{(l)}\in\mathbb{R}^{M_{l-1}\times M_{l}}$ and $\widetilde{\omega}^{(l)}$ is the weight matrices of $\widetilde{f}^{(l)}$ which is filters of the $l$-th layer of $\widetilde{F}^{(L)}$ again with $M_{l}$ neurons. For clarity, we consistently treat weight matrices as high-dimensional vectors. If $\omega^{(l)}\in\mathbb{R}^{M_{l-1}\times M_{l}}$ represents the weight matrix of layer $l$, its vectorized version is denoted by $\omega^{(l)}\in\mathbb{R}^{(M_{l-1}\times M_{l})\times 1}$, where $M_{l}$ is the number of neurons in layer $l$. Note that $\omega^{(l)}$ is referred to as a matrix whenever we use the term "matrix"; otherwise, it is considered as a vector.
@@
In our approach, $\omega^{(l)}$ is projected, using a projection map $\mathcal{P}:\mathbb{R}^{(M_{l-1}\times M_{l})\times 1}\mapsto\mathbb{R}^{(M_{l-1}\times M_{l})\times 1}$, to a probability space $\mathcal{Z}$ with random variables $Z$ and with a distribution, denoted by $P_{\omega^{(l)}}$, which has $\omega^{(l)}$ as its parameter set. Similarly, a probability space $\mathcal{\widetilde{Z}}$ with latent variables $\widetilde{Z}$ and distribution $P_{\widetilde{\omega}^{(l)}}$ with $\widetilde{\omega}^{(l)}$ as its parameter set is assigned to $\widetilde{\omega}^{(l)}$ (we illustrate this in Fig. 1). We show that by computing the Kullback-Leibler (KL) divergence between these two distributions, $KL(P_{\omega^{(l)}}||P_{\widetilde{\omega}^{(l)}})$, we explain the behavior of the pruned network and how magnitude-based pruning maintain the task learning performance. We show that as pruned network’s performance convergences to optimal unpruned network performance, the KL divergence between $P_{\omega^{(l)}}$ and $P_{\widetilde{\omega}^{(l)}}$ for given layer $l$ convergences to zero. Note that what distinguishes our method from Bayesian inference is that the networks considered in this paper are non-Bayesian networks, meaning that there is no prior distribution associated with the network parameters. Hence, the distributions are not a distribution of network parameters but rather they are the distribution of defined latent spaces to clarify the network pruning methods.
@@
Why our approach is important? Our new framework enables a deeper understanding of the complex interplay between network pruning and probabilistic distributions, where the weight tensors serve as their key parameters. Our approach effectively explains the sparsity of networks using latent spaces, shedding light on the interpretability of pruned models.
@@
Notations. For clarity and ease of notation, we define all relevant symbols below. As mentioned earlier, we treat $\omega^{(l)}$ as a weight vector by default unless explicitly mentioned as a matrix using the term "matrix". Consider two neural networks, denoted as $F^{(L)}$ and $\widetilde{F}^{(L)}$ with $L$ layers and filters of their $l$-th layer, i.e., $f^{(l)}$ and $\widetilde{f}^{(l)}$. We use $\omega^{(l)}$ and $\widetilde{\omega}^{(l)}$ to represent the weight matrices of the $l$-th layer in $F^{(L)}$ and $\widetilde{F}^{(L)}$, respectively. Additionally, $\omega^{*}$ and $\widetilde{\omega}^{*}$ as the optimal weight tensors for $F^{(L)}$ and $\widetilde{F}^{(L)}$. The performance difference between these two networks is quantified by the function $D(F,\widetilde{F})$. To measure the dissimilarity between distributions, we use the Kullback–Leibler divergence denoted as $KL(.||.)$ though the intuition behind our theoretical explanations is based on any distance metrix. Finally, the terms $m$, $\sigma$, and $\mathbb{E}$ correspond to a binary mask, activation function, and expectation operation, respectively.
2 Problem Formulation
@@
$\mathbb{E}_{(\mathbf{X},Y)\sim D}\big{\{}L(F^{(L)}(\mathbf{X}),Y)\big{\}}=-\mathbb{E}_{(\mathbf{X},Y)\sim D}\big{\{}Y\cdot\big{(}b+\sum\limits_{F\in\mathcal{F}}\omega_{F}\cdot F^{(L)}(\mathbf{X})\big{)}\big{\}}.$ (1)
@@
Definition 1.
Analogous Projected Patterns (AP2) - Consider the projected map function $\mathcal{P}~{}:~{}\mathbb{R}^{(M_{l-1}\times M_{l})\times 1}\mapsto\mathbb{R}^{(M_{l-1}\times M_{l})\times 1}$. We say filter $f^{(l)}$ and its sparse version $\widetilde{f}^{(l)}$ have AP2 if for small $\epsilon^{(l)}$, $\left\lVert\mathcal{P}\big{(}f^{(l)}\big{)}-\mathcal{P}\big{(}\widetilde{f}^{(l)}\big{)}\right\rVert^{2}\leq\epsilon^{(l)}$. Equivalently, we say layer $l$ with weight $\omega^{(l)}$ and it’s sparse version with sparse weight $\widetilde{\omega}^{(l)}$ have AP2 if for small $\epsilon^{(l)}$, $\left\lVert\mathcal{P}\big{(}\omega^{(l)}\big{)}-\mathcal{P}\big{(}\widetilde{\omega}^{(l)}\big{)}\right\rVert^{2}\leq\epsilon^{(l)}$ or if $\left\lVert(\omega^{(l)}-\widetilde{\omega}^{(l)})\right\rVert^{2}\leq\epsilon^{(l)}$ when we use identity map as the projection function $\mathcal{P}$.
Figure 1: (Left) Proposed latent space $\mathcal{Z}$ and $\widetilde{\mathcal{Z}}$ from Gaussian distribution with mean parameters $\omega$ and $\widetilde{\omega}$ and covariances $\Sigma$ and $\widetilde{\Sigma}$, respectively: $P_{\omega^{(l)}}=G(\omega^{(l)},\Sigma_{1})$ and $P_{\widetilde{\omega}^{(l)}}=G(\widetilde{\omega}^{(l)},\Sigma_{2})$. (Right) A detailed illustration of maping weights matrices to the latent space as the distribution parameter.
Suppose latent variables $Z\in\mathcal{Z}$ and $\widetilde{Z}\in\widetilde{\mathcal{Z}}$ have distributions $P_{\omega}:=P(.,\omega)$ and $P_{\widetilde{\omega}}:=P(.,\widetilde{\omega})$ with parameter sets $\omega$ and $\widetilde{\omega}$, respectively. $\mathcal{Z}$ and $\widetilde{\mathcal{Z}}$ are latent spaces that could overlap. The Kullback-Leibler (KL) divergence [9] between $P_{\omega}$ and $P_{\widetilde{\omega}}$ is
$KL(P_{\omega}\|P_{\widetilde{\omega}}):=\mathbb{E}_{P_{\omega}}\left[\log\left({P_{\omega}}/{P_{\widetilde{\omega}}}\right)\right].$ (2)
Definition 2.
@@
Throughout this paper, we assume that all distributions have positive definite covariance matrix.
Note that the projection map in Definitions 1 and 2 are not necessarily the same. To clarify this, we provide an example that shows if layer $l$ and its sparse version have AP2 with an identity projection map, then they have AP3 property with Gaussian projection i.e. projection $\mathcal{P}$ maps $\omega^{(l)}$ and $\widetilde{\omega}^{(l)}$ to the latent space when latent variables $Z$ and $\widetilde{Z}$ have Gaussian distributions with means $\omega^{(l)}$ and $\widetilde{\omega}^{(l)}$, respectively. Example - Gaussian Projection: Suppose layer $l$ with weight $\omega^{(l)}$ and its sparse version with weight $\widetilde{\omega}^{(l)}$ have AP3 property with Gaussian projection i.e $P_{\omega^{(l)}}\sim G(\omega^{(l)},\Sigma)$ and $P_{\widetilde{\omega}^{(l)}}\sim G(\widetilde{\omega}^{(l)},\widetilde{\Sigma})$, where $\Sigma$ and $\widetilde{\Sigma}$ are covariance matrices. Using (2) we can write
$KL({P}_{\omega^{(l)}}\|P_{\widetilde{\omega}^{(l)}})=\frac{1}{2}\Big{(}\log\frac{|\Sigma|}{|\widetilde{\Sigma}|}+tr(\Sigma^{-1}\widetilde{\Sigma})+(\omega^{(l)}-\widetilde{\omega}^{(l)})^{T}\Sigma^{-1}(\omega^{(l)}-\widetilde{\omega}^{(l)})-d\Big{)},$ (3)
@@
$KL({P}_{\omega^{(l)}}\|P_{\widetilde{\omega}^{(l)}})\leq\frac{1}{2}\left(C+\lambda_{max}\epsilon^{(l)}\right),\;\;\hbox{where $C=\log\frac{|\Sigma|}{|\widetilde{\Sigma}|}+tr(\Sigma^{-1}\widetilde{\Sigma})-d$}$ (4)
@@
$\|\omega^{(l)}-\widetilde{\omega}^{(l)}\|^{2}_{2}\leq\lambda_{min}\|\omega^{(l)}-\widetilde{\omega}^{(l)}\|^{2}_{2}\leq(\omega^{(l)}-\widetilde{\omega}^{(l)})^{T}\Sigma^{-1}(\omega^{(l)}-\widetilde{\omega}^{(l)}))$ (5)
@@
Definition 3.
[10] (Pinsker’s inequality) If $P$ and $Q$ are two distributions on a measurable space $(X,\mathcal{F})$, then $KL(P\|Q)\geq 2d_{TV}^{2}$.
Definition 4.
[9] Consider a measurable space $(\Omega,\mathcal{A})$ and probability measures $P$ and $Q$ defined on $(\Omega,\mathcal{A})$ on a finite domain $D$. The total variation distance between $P$ and $Q$ is defined as
@@
Lemma 1.
[11] Let $P$ and $Q$ be two arbitrary distributions in $\mathbb{R}^{d}$ with mean vectors of $m_{P}$ and $m_{Q}$ and covariance matrices of $\Sigma_{P}$ and $\Sigma_{Q}$. Then, by setting constant $a:=m_{P}-m_{Q}$,
@@
Corollary 1.
Let $P$ and $Q$ be two arbitrary distributions in $\mathbb{R}^{d}$ with mean vectors of $m_{P}$ and $m_{Q}$ and covariance matrices of $\Sigma_{P}$ and $\Sigma_{Q}$, respectively. Set constant $a:=m_{p}-m_{Q}$. Using definition 3 and Lemma 1, $KL(P\|Q)\geq 2d^{2}_{TV}(P,Q)\geq 2\left(\frac{a^{T}a}{2(tr(\Sigma_{P})+tr(\Sigma_{Q}))+a^{T}a}\right)^{2}$.
According to the Corollary 1, for any two continuous distributions $P_{\omega^{(l)}}$ and $P_{\widetilde{\omega}^{(l)}}$ in a $d$-dimensional space, both with identity matrix $I\in\mathbb{R}^{d\times d}$ as their covariance matrices, if AP3 is satisfied, then, AP2 holds when $\epsilon<2$. However, it’s important to note that there can be instances where AP3 doesn’t necessarily imply AP2, and this relationship depends on the specific values of $\epsilon$ and the characteristics of the distributions.
Definition 5.
@@
$|KL(P_{\omega^{*}}\|P_{\widetilde{\omega}^{*}})-KL(P_{\omega^{*}}\|P_{\widetilde{\omega}^{t}})|=O(\epsilon_{t}),\;\;\hbox{where as $\epsilon_{t}\rightarrow 0$ when $t\rightarrow\infty$.}$ (6)
3 Theoretical Analysis
In DNNs compression, it is essential to maintain the performance with a core mission of reducing computational complexity and memory usage. However, in many compression methods, after significant computational improvement, we want the performance of the network after compression to be as close as possible to the best performance of the complete (non-compressed) network.
Definition 6.

### v3

@@ Abstract.
Despite the impressive performance of deep neural networks (DNNs), their computational complexity and storage space consumption have led to the concept of network compression. While DNN compression techniques such as pruning and low-rank decomposition have been extensively studied, there has been insufficient attention paid to their theoretical explanation. In this paper, we propose a novel theoretical framework that leverages a probabilistic latent space of DNN weights and explains the optimal network sparsity by using the information-theoretic divergence measures. We introduce new analogous projected patterns (AP2) and analogous-in-probability projected patterns (AP3) notions for DNNs and prove that there exists a relationship between AP3/AP2 property of layers in the network and its performance. Further, we provide a theoretical analysis that explains the training process of the compressed network. The theoretical results are empirically validated through experiments conducted on standard pre-trained benchmarks, including AlexNet, ResNet50, VGG16, and ViT using CIFAR10, CIFAR100, and Tiny-ImageNet datasets. Through our experiments, we highlight the relationship of AP3 and AP2 properties with fine-tuning pruned DNNs and sparsity levels.
Network Pruning, Theoritical Explanation, Latent Space
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: TOPMLjournalvolume: 00journalnumber: 0article: 1publicationmonth: 0ccs: Mathematics of computing Probability and statisticsccs: Computing methodologies Machine learning
1. Introduction
@@
A wide variety of real-world problems are investigated by deep neural networks (DNNs) which enable powerful learning across a wide range of task learning. DNNs have provided remarkable performance in solving complex problems and this has led to significant growth of interest in the area of neural networks. However; the higher the performance, the more the number of parameters, resulting in complexity and space consumption. While DNNs pruning methods (Yeom et al., 2021; Blalock et al., 2020; Ding et al., 2019a; Lee et al., 2020; Park et al., 2020; Ding et al., 2019b; Liu et al., 2021; You et al., 2019) compress networks and tackle computational complexity challenges, due to the lack of sufficient theoretical explanation of such methods, it’s difficult to explain their behaviors. In this paper, we propose a novel framework that explains DNN pruning using probabilistic latent spaces of deep network’s weights and divergence measures (Sason, 2022; Chen and Frid, 2001; Pradier et al., 2018). Our main idea in this work is built upon mapping the weight matrix of the layers in both the original network and its corresponding sparse/pruned network to probability spaces from a distribution where the weight matrices are the parameter sets. To clarify our approach, consider a network $F^{(L)}$ with $L$-layers and its sparse version, denoted by $\widetilde{F}^{(L)}$ with parameters $\omega$ and $\widetilde{\omega}=m\odot\omega$, respectively, in which $m$ is a binary mask. Suppose that $f^{(l)}$, filters of the $l$-th layer of $F^{(L)}$ with $M_{l}$ neurons has weight matrix of $\omega^{(l)}\in\mathbb{R}^{M_{l-1}\times M_{l}}$ and $\widetilde{\omega}^{(l)}$ is the weight matrices of $\widetilde{f}^{(l)}$ which is filters of the $l$-th layer of $\widetilde{F}^{(L)}$ again with $M_{l}$ neurons. For clarity, we consistently treat weight matrices as high-dimensional vectors. If $\omega^{(l)}\in\mathbb{R}^{M_{l-1}\times M_{l}}$ represents the weight matrix of layer $l$, its vectorized version is denoted by $\omega^{(l)}\in\mathbb{R}^{(M_{l-1}\times M_{l})\times 1}$, where $M_{l}$ is the number of neurons in layer $l$. Note that $\omega^{(l)}$ is referred to as a matrix whenever we use the term ”matrix”; otherwise, it is considered as a vector.
@@
In our approach, $\omega^{(l)}$ is projected, using a projection map $\mathcal{P}:\mathbb{R}^{(M_{l-1}\times M_{l})\times 1}\mapsto\mathbb{R}^{(M_{l-1}\times M_{l})\times 1}$, to a probability space $\mathcal{Z}$ with random variables $Z$ and with a distribution, denoted by $P_{\omega^{(l)}}$, which has $\omega^{(l)}$ as its parameter set. Similarly, a probability space $\mathcal{\widetilde{Z}}$ with latent variables $\widetilde{Z}$ and distribution $P_{\widetilde{\omega}^{(l)}}$ with $\widetilde{\omega}^{(l)}$ as its parameter set is assigned to $\widetilde{\omega}^{(l)}$ (we illustrate this in Fig. 1). We show that by computing the Kullback-Leibler (KL) divergence between these two distributions, $KL(P_{\omega^{(l)}}||P_{\widetilde{\omega}^{(l)}})$, we explain the behavior of the pruned network and how deep network pruning maintains the task learning performance. We show that as the pruned network’s performance convergences to optimal unpruned network performance, the KL divergence between $P_{\omega^{(l)}}$ and $P_{\widetilde{\omega}^{(l)}}$ for given layer $l$ convergences to zero. Note that what distinguishes our method from Bayesian inference is that the networks considered in this paper are non-Bayesian networks, meaning that there is no prior distribution associated with the network parameters. Hence, the distributions are not a distribution of network parameters but rather they are the distribution of defined latent spaces to clarify the network pruning methods.
Core idea of our analysis: For given neural network $F^{(L)}$ with $L$ layers, the parameters of layer $l$, $\omega^{(l)}$ are sparsified (pruned) and the weights of the layer $l$ after pruning is $\widetilde{\omega}^{(l)}$. Both $\omega^{(l)}$ and $\widetilde{\omega}^{(l)}$ are the parameters of a latent distribution with variables $Z$ and $\widetilde{Z}$. Our approach involves projecting the network’s weights—both before and after pruning—into a probabilistic latent space with different distributions, such as Gaussian. The mapping (projection) is equivalent to assuming the latent variables $Z^{(l)}$ and $\widetilde{Z}^{(l)}$ have distributions $P_{\omega^{(l)}}$ and $P_{\widetilde{\omega}^{(l)}}$, respectively. Next, by computing the distance between distributions $P_{\omega^{(l)}}$ and $P_{\widetilde{\omega}^{(l)}}$, we monitor and explain the behavior of magnitude based pruning. In this work we apply KL divergence, $KL(P_{\omega^{(l)}}||P_{\widetilde{\omega}^{(l)}})$ as the distance between the distributions. In addition, we consider the distributions $P_{\omega^{(l)}}$ and $P_{\widetilde{\omega}^{(l)}}$ are uniform and the distacne is $L_{2}$-norm, $\|\omega^{(l)}-\widetilde{\omega}^{(l)}\|_{2}^{2}$.
Note that, in our analysis, this projection is applied to the weights of the network, therefore all weight-based pruning techniques can benefit from our analytical investigations. Further, because other pruning techniques impact the weights of the network directly or indirectly, they benefit from our analytical investigations in this paper as well.
Structure of the projection operator: To analyze the structure of the proposed projection, suppose two deep networks $F^{(L)}_{1}$ and $F^{(L)}_{2}$ with $\omega_{1}$ and $\omega_{2}$ weight matrices respectively. While training $F^{(L)}_{1}$ and $F^{(L)}_{2}$ if the average of $\omega_{1}$ and $\omega_{2}$ weights over all training sample convergence to two significantly different values, with high probability their performance difference increases. Now suppose the weights are the mean parameter of a $d$-dimensional Gaussian distribution (or generally weights are the parameters of a given distribution $P_{\omega}$). This is equivalent to projecting the weights into a latent space $Z$ with distribution $P_{\omega}$. To generalize our projection, we consider the KL divergence between distributions. In this work, $\omega_{1}=\omega$ (original weights) and $\omega_{2}=\widetilde{\omega}$ (prunned weights). Additional analysis has been provided in Section 3 - Theoretical Analysis Section.
@@
Why our explainable approach is important? Our new framework enables a deeper understanding of the complex interplay between network pruning and probabilistic distributions, where the weight tensors serve as their key parameters. Our approach effectively explains the sparsity of networks using latent spaces, shedding light on the interpretability of pruned models.
In this work, we focus on weight pruning for the key reasons below:
  • Relevance and Familiarity: Weight pruning is one of the most commonly used pruning techniques in neural network compression, making it a logical choice for our experiments. By focusing on this method, we compare our findings with a wide body of existing research, helping to place our work in context.
  • Precision in Optimization: Since our paper is about achieving optimal pruning, weight pruning is particularly useful. It allows for fine-tuning of the network’s sparsity, helping us strike a balance between maintaining accuracy and improving efficiency. This kind of precision is harder to achieve with methods like structural pruning, which can make broader, less targeted changes to the network.
  • Efficiency: Weight pruning is often less computationally intensive compared to other pruning methods, especially when dealing with overparametrized networks. The magnitude-based pruning techniques it practical for cutting-edge practical applications due to its efficiency. In our experiments, we thoroughly test our theoretical framework across various conditions.
@@
Notations. For clarity and ease of notation, we define all relevant symbols below. As mentioned earlier, we treat $\omega^{(l)}$ as a weight vector by default unless explicitly mentioned as a matrix using the term ”matrix”. Consider two neural networks, denoted as $F^{(L)}$ and $\widetilde{F}^{(L)}$ with $L$ layers and filters of their $l$-th layer, i.e., $f^{(l)}$ and $\widetilde{f}^{(l)}$. We use $\omega^{(l)}$ and $\widetilde{\omega}^{(l)}$ to represent the weight matrices of the $l$-th layer in $F^{(L)}$ and $\widetilde{F}^{(L)}$, respectively. Additionally, $\omega^{*}$ and $\widetilde{\omega}^{*}$ as the optimal weight tensors for $F^{(L)}$ and $\widetilde{F}^{(L)}$. The performance difference between these two networks is quantified by the function $D(F,\widetilde{F})$. To measure the dissimilarity between distributions, we use the Kullback–Leibler divergence denoted as $KL(.||.)$ though the intuition behind our theoretical explanations is based on any distance metric. Finally, the terms $m$, $\sigma$, and $\mathbb{E}$ correspond to a binary mask, activation function, and expectation operation, respectively.
2. Problem Formulation
@@
(1) $\mathbb{E}_{(\mathbf{X},Y)\sim D}\big{\{}L(F^{(L)}(\mathbf{X}),Y)\big{\}}=-\mathbb{E}_{(\mathbf{X},Y)\sim D}\big{\{}Y\cdot\big{(}b+\sum\limits_{F\in\mathcal{F}}\omega_{F}\cdot F^{(L)}(\mathbf{X})\big{)}\big{\}}.$
@@
Definition 2.1.
Analogous Projected Patterns (AP2) - Consider the projected map function $\mathcal{P}:\mathbb{R}^{(M_{l-1}\times M_{l})\times 1}\mapsto\mathbb{R}^{(M_{l-1}\times M_{l})\times 1}$. We say filter $f^{(l)}$ and its sparse version $\widetilde{f}^{(l)}$ have AP2 if for small $\epsilon^{(l)}$, $\left\lVert\mathcal{P}\big{(}f^{(l)}\big{)}-\mathcal{P}\big{(}\widetilde{f}^{(l)}\big{)}\right\rVert^{2}\leq\epsilon^{(l)}$. Equivalently, we say layer $l$ with weight $\omega^{(l)}$ and it’s sparse version with sparse weight $\widetilde{\omega}^{(l)}$ have AP2 if for small $\epsilon^{(l)}$, $\left\lVert\mathcal{P}\big{(}\omega^{(l)}\big{)}-\mathcal{P}\big{(}\widetilde{\omega}^{(l)}\big{)}\right\rVert^{2}\leq\epsilon^{(l)}$ or if $\left\lVert(\omega^{(l)}-\widetilde{\omega}^{(l)})\right\rVert^{2}\leq\epsilon^{(l)}$ when we use identity map as the projection function $\mathcal{P}$.
Figure 1. (Left) Proposed latent space $\mathcal{Z}$ and $\widetilde{\mathcal{Z}}$ from Gaussian distribution with mean parameters $\omega$ and $\widetilde{\omega}$ and covariances $\Sigma$ and $\widetilde{\Sigma}$, respectively: $P_{\omega^{(l)}}=G(\omega^{(l)},\Sigma_{1})$ and $P_{\widetilde{\omega}^{(l)}}=G(\widetilde{\omega}^{(l)},\Sigma_{2})$. (Right) A detailed illustration of mapping weights matrices to the latent space as the distribution parameter.
Suppose latent variables $Z\in\mathcal{Z}$ and $\widetilde{Z}\in\widetilde{\mathcal{Z}}$ have distributions $P_{\omega}:=P(.,\omega)$ and $P_{\widetilde{\omega}}:=P(.,\widetilde{\omega})$ with parameter sets $\omega$ and $\widetilde{\omega}$, respectively. $\mathcal{Z}$ and $\widetilde{\mathcal{Z}}$ are latent spaces that could overlap. The Kullback-Leibler (KL) divergence (Cover and Thomas, 1991) between $P_{\omega}$ and $P_{\widetilde{\omega}}$ is
(2) $KL(P_{\omega}\|P_{\widetilde{\omega}}):=\mathbb{E}_{P_{\omega}}\left[\log\left({P_{\omega}}/{P_{\widetilde{\omega}}}\right)\right].$
Definition 2.2.
@@
Throughout this paper, we assume that all distributions have a positive definite covariance matrix, defined as follows. For the diagonal covariance matrix, we use $\Sigma=\sigma^{2}I$, where $\sigma>0$ is a positive scalar to ensure that the matrix is positive definite, and $I$ is the identity matrix, which ensures that the covariance matrix is diagonal. For the non-diagonal covariance matrix, we use $\Sigma=\beta V\Lambda V^{T}$, where $\beta>0$ is a scalar used to control the scale of the covariance matrix, and $V$ is the eigenvector matrix derived from $R+R^{T}$, where $R$ is a random matrix. This ensures the covariance matrix is non-diagonal and symmetric, and $\Lambda=(\lambda_{1},\ldots,\lambda_{n})$ is the diagonal matrix of eigenvalues, where $n$ is the dimension of the covariance matrix and $\lambda_{i}>0$. This guarantees that the matrix is positive definite.
Note that the projection map in Definitions 2.1 and 2.2 are not necessarily the same. To clarify this, we provide an example that shows if layer $l$ and its sparse version have AP2 with an identity projection map, then they have AP3 property with Gaussian projection i.e. projection $\mathcal{P}$ maps $\omega^{(l)}$ and $\widetilde{\omega}^{(l)}$ to the latent space when latent variables $Z$ and $\widetilde{Z}$ have Gaussian distributions with means $\omega^{(l)}$ and $\widetilde{\omega}^{(l)}$, respectively.
Example - Gaussian Projection: Suppose layer $l$ with weight $\omega^{(l)}$ and its sparse version with weight $\widetilde{\omega}^{(l)}$ have AP3 property with Gaussian projection i.e $P_{\omega^{(l)}}\sim G(\omega^{(l)},\Sigma)$ and $P_{\widetilde{\omega}^{(l)}}\sim G(\widetilde{\omega}^{(l)},\widetilde{\Sigma})$, where $\Sigma$ and $\widetilde{\Sigma}$ are covariance matrices. Using (2) we can write
(3) $KL({P}_{\omega^{(l)}}\|P_{\widetilde{\omega}^{(l)}})=\frac{1}{2}\Big{(}\log\frac{|\Sigma|}{|\widetilde{\Sigma}|}+tr(\Sigma^{-1}\widetilde{\Sigma})+(\omega^{(l)}-\widetilde{\omega}^{(l)})^{T}\Sigma^{-1}(\omega^{(l)}-\widetilde{\omega}^{(l)})-d\Big{)},$
@@
(4) $KL({P}_{\omega^{(l)}}\|P_{\widetilde{\omega}^{(l)}})\leq\frac{1}{2}\left(C+\lambda_{max}\epsilon^{(l)}\right),\;\;\hbox{where $C=\log\frac{|\Sigma|}{|\widetilde{\Sigma}|}+tr(\Sigma^{-1}\widetilde{\Sigma})-d$.}$
@@
(5) $\|\omega^{(l)}-\widetilde{\omega}^{(l)}\|^{2}_{2}\leq\lambda_{min}\|\omega^{(l)}-\widetilde{\omega}^{(l)}\|^{2}_{2}\leq(\omega^{(l)}-\widetilde{\omega}^{(l)})^{T}\Sigma^{-1}(\omega^{(l)}-\widetilde{\omega}^{(l)})).$
@@
Definition 2.3.
(Csiszár and Körner, 2011) (Pinsker’s inequality) If $P$ and $Q$ are two distributions on a measurable space $(X,\mathcal{F})$, then $KL(P\|Q)\geq 2d_{TV}^{2}$.
Definition 2.4.
(Cover and Thomas, 1991) Consider a measurable space $(\Omega,\mathcal{A})$ and probability measures $P$ and $Q$ defined on $(\Omega,\mathcal{A})$ on a finite domain $D$. The total variation distance between $P$ and $Q$ is defined as
@@
Lemma 2.5.
(Nishiyama, 2022) Let $P$ and $Q$ be two arbitrary distributions in $\mathbb{R}^{d}$ with mean vectors of $m_{P}$ and $m_{Q}$ and covariance matrices of $\Sigma_{P}$ and $\Sigma_{Q}$. Then, by setting constant $a:=m_{P}-m_{Q}$,
@@
Corollary 2.6.
Let $P$ and $Q$ be two arbitrary distributions in $\mathbb{R}^{d}$ with mean vectors of $m_{P}$ and $m_{Q}$ and covariance matrices of $\Sigma_{P}$ and $\Sigma_{Q}$, respectively. Set constant $a:=m_{p}-m_{Q}$. Using definition 2.3 and Lemma 2.5, $KL(P\|Q)\geq 2d^{2}_{TV}(P,Q)\geq 2\left(\frac{a^{T}a}{2(tr(\Sigma_{P})+tr(\Sigma_{Q}))+a^{T}a}\right)^{2}$.
According to the Corollary 2.6, for any two continuous distributions $P_{\omega^{(l)}}$ and $P_{\widetilde{\omega}^{(l)}}$ in a $d$-dimensional space, both with identity matrix $I\in\mathbb{R}^{d\times d}$ as their covariance matrices, if AP3 is satisfied, then, AP2 holds when $\epsilon<2$. However, it’s important to note that there can be instances where AP3 doesn’t necessarily imply AP2, and this relationship depends on the specific values of $\epsilon$ and the characteristics of the distributions.
Definition 2.7.
@@
(6) $|KL(P_{\omega^{*}}\|P_{\widetilde{\omega}^{*}})-KL(P_{\omega^{*}}\|P_{\widetilde{\omega}^{t}})|=O(\epsilon_{t}),\;\;\hbox{where as $\epsilon_{t}\rightarrow 0$ when $t\rightarrow\infty$.}$
Definition 2.8.